OpenStudy (anonymous):

A garden is divided into 12 sections of equal size. In 5 of these sections you plant tomatoes. an armadillo dug up 1/3 of these plants. What fraction of the garden now contains tomatoe plants? Please help...

OpenStudy (anonymous):

\[\frac{ 5 }{ 12 }-\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

I did that exact problem. And got the answer of 1/12. But my teacher said the correct answer is 5/18. Im confused on how she got that fraction.

OpenStudy (anonymous):

\[\frac{ 5 }{ 3 }\]dug up from the tomatoes \[\frac{ \frac{ 5 }{ 3 } }{ 12}=\frac{ 5 }{ 36 }\] dug up from the whole garden \[\frac{ 5 }{ 12 }\] the five sections containing tomatoes \[\frac{ 5 }{ 12 }-\frac{ 5 }{ 36 }=\frac{ 5 }{ 18 }\]

OpenStudy (anonymous):

Ohhh.... thanks so much!!!!